meganmoney543:

John was asked to solve the following system using multiplication and addition. He tried twice to come up with an equivalent set of equations. Analyze each of John's attempts. Tell whether the attempt is an accurate and equivalent system of equations. Describe, in complete sentences, which operations were completed in each system and tell why or why not they were accurate operations. Original System of Equations: 7x - 3y = 4 2x - 4y = 1 Attempt #1: 28x - 12y = 16 -6x + 12y = -3 Attempt #2: 14x - 6y = 4 -14x + 28y = 1

Rigo:

fist..whut do u think..?

meganmoney543:

i used a calculator for the 2 attempts and it dont got the right answer

meganmoney543:

can u use basketball or baseball terms when u explain plz

Rigo:

the u=numbers r x and y..u cant multiply those

meganmoney543:

ok

Rigo:

the second attempt was the better of the rwo

meganmoney543:

ok

meganmoney543:

do you know what methods he used to solve the problems tho

meganmoney543:

@new2luv

new2luv:

what do you have to do here find the answer

snowflake0531:

Do you see that he tried to get at least one variable equal? this means he's using the elimination method

meganmoney543:

yes

snowflake0531:

His first attempt was to cancel out the y's, so he multiplied the equations by 4 and -3

snowflake0531:

In this way, when he cancels, he will be left with x

meganmoney543:

yes is there a way you can use basketball or sports terms if not it is ok

snowflake0531:

7x - 3y = 4 2x - 4y = 1 Attempt #1: 28x - 12y = 16 -6x + 12y = -3 Let's first analyze the first one 7x-3y=4 28x-12y=16 Do you think he multiplied it correctly? (he multiplied the equation by 4)

meganmoney543:

and no he didnt

snowflake0531:

what

snowflake0531:

Why not? where is his mistake then

meganmoney543:

hol up lemme think

snowflake0531:

your thoughts now...?

meganmoney543:

im working it out on paper but i think that he multiplied it wrong bc the answer wasnt correct on the 2nd problem in the first attempt